OpenStudy (anonymous):

the product of 2 positive intergers is 90.find the 2 intergers if the larger is 3 more than twice the smaller

OpenStudy (anonymous):

say, one integer is x the other one is 2x+3 (as per the given condition) \[x(2x+3)=90\] or, \[2x ^{2}+3x-90=0\] or,\[(2x+15)(x-6)=0\] so, x is either -15/2 or 6 but x is a positive integer (given) so, x=6